\begin{thebibliography}{10}

\bibitem{QMBlokhintsev}
D.~I. Blokhintsev, {\em Quantum Mechanics}.
\newblock Dordrecht: D. Reidel Publishing Company, 1964.
\newblock Translated from the Russian edition.

\bibitem{QMDavydov}
A.~S. Davydov, {\em Quantum Mechanics}.
\newblock Oxford: Pergamon Press, 2nd~ed., 1976.

\bibitem{Me24NCevo}
A.~P. Meilakhs, ``Non-coherent evolution of closed weakly interacting system
  leads to equidistribution of probabilities of microstates,'' {\em arXiv
  preprint}, 2024.

\bibitem{Me24transmission}
A.~Meilakhs, ``Transmission of waves and particles through the interface:
  reversibility and coherence,'' {\em Annals of Physics}, vol.~466, p.~169686,
  2024.

\bibitem{Me25Larotonda}
A.~P. Meilakhs, C.~Pastorino, and M.~Larotonda, ``Irreversibility of heat
  transport and optical non-coherence: an experimental demonstration of the
  analogy,'' {\em arXiv preprint}, 2025.
\newblock Version v2 (submitted 14 May 2025, revised 16 Sept 2025).

\bibitem{Joos2003appearance}
E.~Joos, ed., {\em Decoherence and the appearance of a classical world in
  quantum theory}.
\newblock Physics and astronomy online library, Berlin: Springer, 2. ed.~ed.,
  2003.
\newblock Hier auch später erschienene, unveränd. Nachdr.

\bibitem{Zurek1993}
W.~H. Zurek, ``Preferred states, predictability, classicality and the
  environment-induced decoherence,'' {\em Progress of Theoretical Physics},
  vol.~89, pp.~281--312, 1993.

\bibitem{Zurek1981packet}
W.~H. Zurek, ``Pointer basis of quantum apparatus: Into what mixture does the
  wave packet collapse?,'' {\em Physical Review D}, vol.~24, pp.~1516--1525,
  1981.

\bibitem{Joos1985packet}
E.~Joos and H.~D. Zeh, ``The emergence of classical properties through
  interaction with the environment,'' {\em Zeitschrift f{\"u}r Physik B
  Condensed Matter}, vol.~59, pp.~223--243, 1985.

\bibitem{Omnes2011packet}
R.~Omn\`es, ``Decoherence and wave function collapse,'' {\em Foundations of
  Physics}, vol.~41, no.~12, pp.~1857--1880, 2011.

\bibitem{Zurek2013packet}
W.~H. Zurek, ``Wave-packet collapse and the core quantum postulates:
  Discreteness of quantum jumps from unitarity, repeatability, and actionable
  information,'' {\em Physical Review A}, vol.~87, no.~5, p.~052111, 2013.

\bibitem{Pessoa1997contr}
J.~O.~Pessoa, ``Can the decoherence approach help to solve the measurement
  problem?,'' {\em Synthese}, vol.~113, pp.~323--346, 1997.

\bibitem{Adler2003contr}
S.~L. Adler, ``Why decoherence has not solved the measurement problem: A
  response to p. w. anderson,'' {\em Studies in History and Philosophy of
  Modern Physics}, vol.~34, pp.~135--142, 2003.

\bibitem{Wallace2012contr}
D.~Wallace, ``The emergence of classicality from quantum theory,'' {\em
  Philosophical Transactions of the Royal Society A: Mathematical, Physical and
  Engineering Sciences}, vol.~370, pp.~4576--4593, 2012.

\bibitem{Harris2004}
S.~Harris, {\em An introduction to the theory of the Boltzmann equation}.
\newblock Dover books on physics, Mineola, N.Y. [u.a.]: Dover Publications,
  2004.
\newblock Unabridged republ. of the work originally publ. by Holt, Rinehart and
  Winston, New York, 1971.

\bibitem{Grad1958}
H.~Grad, ``Principles of the kinetic theory of gases,'' in {\em Handbuch der
  Physik}, vol.~12, Berlin: Springer, 1958.

\bibitem{Lanford1975}
O.~E. Lanford, ``Time evolution of large classical systems,'' in {\em Lecture
  Notes in Physics} (J.~Moser, ed.), vol.~38, pp.~1--111, Berlin:
  Springer-Verlag, 1975.

\bibitem{Cohen1960InCondAsymm}
E.~Cohen and T.~Berlin, ``Note on the derivation of the boltzmann equation from
  the liouville equation,'' {\em Physica}, vol.~26, pp.~717--729, Sept. 1960.

\bibitem{Marklof2011Refine}
J.~Marklof and A.~Str{\"o}mbergsson, ``The boltzmann--grad limit of the
  periodic lorentz gas,'' {\em Annals of Mathematics}, vol.~174, pp.~225--298,
  2011.

\bibitem{Mischler2013Refine}
S.~Mischler and C.~Mouhot, ``Kac's program in kinetic theory,'' {\em
  Inventiones Mathematicae}, vol.~193, no.~1, pp.~1--147, 2013.

\bibitem{Bodineau2017Refine}
T.~Bodineau, I.~Gallagher, and L.~Saint-Raymond, ``From hard sphere dynamics to
  the stokes--fourier equations: an {$L^2$} analysis of the boltzmann--grad
  limit,'' {\em Annals of PDE}, vol.~3, 2017.

\bibitem{Deng2024Hil6}
Y.~Deng, Z.~Hani, and X.~Ma, ``Long time derivation of the boltzmann equation
  from hard sphere dynamics,'' {\em arXiv preprint}, 2024.

\bibitem{Deng2025Hil6}
Y.~Deng, Z.~Hani, and X.~Ma, ``Hilbert’s sixth problem: Derivation of fluid
  equations via boltzmann’s kinetic theory,'' {\em arXiv preprint}, 2025.

\bibitem{ArnoldMech}
V.~I. Arnold, {\em Mathematical Methods of Classical Mechanics}, vol.~60 of
  {\em Graduate Texts in Mathematics}.
\newblock New York: Springer, 2nd~ed., 1989.

\bibitem{Zemanian2011}
A.~H. Zemanian, {\em Distribution Theory and Transform Analysis: An
  Introduction to Generalized Functions, with Applications}.
\newblock Courier Corporation, 2011.

\bibitem{ArnoldODE}
V.~I. Arnold, {\em Ordinary Differential Equations}.
\newblock Universitext, Berlin; Heidelberg: Springer, 1992.

\bibitem{Ziman1960}
J.~M. Ziman, {\em Electrons and Phonons: The Theory of Transport Phenomena in
  Solids}.
\newblock Oxford: Oxford University Press, 1960.

\bibitem{LandauV10}
L.~D. Landau and E.~M. Lifshitz, {\em Physical Kinetics}, vol.~10 of {\em
  Course of Theoretical Physics}.
\newblock Oxford: Pergamon Press, 1981.

\bibitem{Anselm1981}
A.~I. Anselm, {\em Introduction to Semiconductor Theory}.
\newblock Englewood Cliffs, NJ: Prentice--Hall, 1981.

\bibitem{breuer2002openquant}
H.-P. Breuer and F.~Petruccione, {\em The theory of open quantum systems}.
\newblock OUP Oxford, 2002.

\bibitem{lindblad1976generators}
G.~Lindblad, ``On the generators of quantum dynamical semigroups,'' {\em
  Communications in mathematical physics}, vol.~48, no.~2, pp.~119--130, 1976.

\bibitem{rigol2008thermalization}
M.~Rigol, V.~Dunjko, and M.~Olshanii, ``Thermalization and its mechanism for
  generic isolated quantum systems,'' {\em Nature}, vol.~452, no.~7189,
  pp.~854--858, 2008.

\bibitem{deutsch2010thermodynamic}
J.~M. Deutsch, ``Thermodynamic entropy of a many-body energy eigenstate,'' {\em
  New Journal of Physics}, vol.~12, no.~7, p.~075021, 2010.

\bibitem{dymarsky2018subsystem}
A.~Dymarsky, N.~Lashkari, and H.~Liu, ``Subsystem eigenstate thermalization
  hypothesis,'' {\em Physical Review E}, vol.~97, no.~1, p.~012140, 2018.

\end{thebibliography}

\documentclass[reprint,secnumarabic,amssymb, nobibnotes, nofootinbib, aps, prd, onecolumn]{revtex4-2}
\usepackage[utf8]{inputenc}
\usepackage{xcolor}
\usepackage{float}
\usepackage{amsmath}
\usepackage{graphicx}

\usepackage{babel}

\begin{document}

\title{The derivation of the Liouville equation from the Schrödinger equation and its implications}

\author{A.\,P. Meilakhs} 
\email[A.\,P. Meilakhs: ]{iestfi@gmail.com}
\affiliation{$^1$Centro Atomico Constituyentes, CNEA, Av. Gral. Paz 1499, San Martin, Buenos Aires, 1650, Argentina }
\date{\today}

\begin{abstract}
We present a new way of deriving classical mechanics from quantum mechanics. A key feature of the method is its compatibility with the standard approach used to derive transition rates between quantum states due to interactions. We apply the developed method to derive the main formulas of physical kinetics. We observe that, through the Liouville equation, we can deduce the non-collision part of the Boltzmann equation, and that, through the matrix of transition rates, we can deduce the collision integral. As a final result of the manuscript, we derive the Boltzmann equation from the Schrödinger equation as a single piece of formal mathematical manipulation, without any non-rigorous plausible reasoning used to glue together its different parts.
\end{abstract}

\maketitle

\section{Introduction}

When deriving classical mechanics from quantum mechanics, standard textbooks on the subject \cite{QMBlokhintsev, QMDavydov} present one of three complementary types of reasoning. One is the quasiclassical approximation, which consists of substituting the wave function in the form $\psi = \exp(iS/\hbar)$ into the Schrödinger equation and expanding it in a power series in $\hbar$. The zeroth term is the only one that survives in the limit $\hbar \to 0$, and it corresponds to the Hamilton–Jacobi equation of classical mechanics. Another is the Ehrenfest theorem, which provides relations between the coordinate and momentum operators analogous to the Hamiltonian equations of classical mechanics.

The last method considers a wave packet and the properties of its motion. It is shown that, under sufficiently good conditions, the probability density moves with velocity $\vec p/m$, and its acceleration is proportional to the gradient of the potential, just as in classical mechanics.

These methods, while they serve the purpose of presenting the transition from quantum to classical almost perfectly, still lack a small element each. The classical Hamilton–Jacobi equation defines a trajectory along which a classical particle actually moves, and along which the change of coordinates occurs. By contrast, the quasiclassical wave function is constant in time, up to a phase factor, so no redistribution of probability takes place. In the same spirit, Ehrenfest’s theorem is too general: it applies to any state of the system, even to stationary states of discrete spectrum, which look nothing like motion along trajectories in classical mechanics.

The wave packet is the concept that most closely resembles motion along a trajectory. However, it lacks generality. What happens when a wave packet encounters a potential barrier and is partly transmitted and partly reflected? If we assume that the potential is sufficiently smooth both before and after the barrier, it is intuitively clear that each part of the wave packet behaves like a classical particle. On the one hand, we can no longer call the state of the system classical, since it now consists of two parts, something that does not occur for a classical particle. On the other hand, the formal derivation of the proportionality between acceleration and the gradient of the potential uses a Taylor expansion of the potential in a small region in which the wave packet is localized. We cannot perform this expansion if the wave function is a superposition of several wave packets that are at a sufficient distance from each other.

Moreover, what if the system consists of many particles, and they are bosonic or fermionic, with the corresponding symmetry requirements for the wave function? Will this affect the behavior of the many-dimensional wave packet? What if the system has interactions that result in quantum transitions, so that probability is continuously redistributed between different wave packets? We want an extension of the wave-packet theory that is able to address these questions.

In the current manuscript, we present a new derivation of classical mechanics from quantum mechanics. More precisely, we derive the Liouville equation from the Schrödinger equation. This derivation has the same generality as the quasiclassical derivation and also captures the actual motion along a trajectory, as in the wave-packet picture.

The one-dimensional case is treated in Section II, and the generalization to an arbitrary number of dimensions is given in Section III. In our previous works, we developed a method for deriving the Boltzmann collision integral from quantum mechanics \cite{Me24NCevo}, based on the ideas of our earlier articles \cite{Me24transmission, Me25Larotonda}. In Section IV of the present manuscript, we combine the results obtained here with those earlier results to derive the full Boltzmann equation from quantum mechanics.

We now want to show how the results of this manuscript are connected to other modern research directions. We can identify two such connections. The first is the derivation of classical mechanics through the concepts of decoherence theory \cite{Joos2003appearance}. As we have already pointed out, a wave packet does not fully represent the behavior of classical particles, since it can evolve into a superposition of states with different probabilities, something that does not occur for classical particles. Decoherence theory attempts to overcome this problem by introducing interactions with the environment as a mechanism for the collapse of the wave function into one particular state \cite{Zurek1993}. Several different approaches exist \cite{Zurek1981packet, Joos1985packet, Omnes2011packet, Zurek2013packet}, along with a well-known set of objections \cite{Pessoa1997contr, Adler2003contr, Wallace2012contr}. The major objection is that decoherence explains only the vanishing of the off-diagonal elements of the density matrix, but does not explain why a single definite outcome is realized in a measurement.

The goal of the theory we present is to deduce the Boltzmann equation from the Schrödinger equation. For this, we only need to explain the classical-like behavior of each sub–wave packet that represents a state; we do not need the exact outcome or a collapse of the state to a single wave packet. Hence we do not encounter the outcome problem.

The second line of research is the derivation of the Boltzmann equation from classical mechanics. The classical approach involves the introduction of the BBGKY hierarchy \cite{Harris2004}, the application of Grad’s limit \cite{Grad1958}, and the subsequent Lanford derivation \cite{Lanford1975}. However, this leads to a contradiction: classical mechanics is time-reversible, whereas the Boltzmann equation is irreversible \cite{Cohen1960InCondAsymm}. As in Boltzmann’s original derivation, one must postulate molecular chaos as a model for the two-particle distribution function, without fully elucidating its origin. The derivation has been refined several times \cite{Marklof2011Refine, Mischler2013Refine, Bodineau2017Refine}. As a result, it has been rigorously shown that the evolution of the distribution function of hard spheres on short timescales, as described by classical mechanics, resembles the description given by the Boltzmann equation.

Very recently, much attention has been drawn to the papers \cite{Deng2024Hil6, Deng2025Hil6}, which claim a proof valid for an arbitrarily long time interval. However, it is not clear how this could overcome the well-known difficulty posed by Poincaré’s recurrence theorem \cite{ArnoldMech}. A system governed by classical mechanics should return arbitrarily close to its initial state (which may be highly nonequilibrium), perhaps after a very long time, whereas the Boltzmann equation predicts that the system reaches a steady state once the relaxation time has passed.

In our theory, we resolve the contradiction between the reversibility of classical mechanics and the irreversibility of the Boltzmann equation in a completely different way. In our previous paper \cite{Me24NCevo}, we proposed a derivation of the collision-integral part of the Boltzmann equation from quantum mechanics. We specifically addressed the problem of the origin of irreversibility: it arises from the finite spectral width of quantum states. However, we note that this theory is applicable only for interactions with sufficiently high-frequency fields. In the present paper, we describe the interaction of the system with sufficiently low-frequency fields. Thus, the non-collision part, which is equivalent to classical mechanics, and the collision part of the Boltzmann equation correspond to two different limits of quantum mechanics -- the long-wavelength and short-wavelength limits, respectively. Hence, there is no contradiction.

\section{One-dimensional single-particle equation}

We start with the one-dimensional Schrödinger equation for a single particle. We seek the spatial part of the wave function in the form
\begin{equation}
\Psi (x) = \sum_p A_p(x) \phi_p(x).
\label{Psi}
\end{equation}
 Here $A_p(x)$ is a smooth envelope of wave function that determines the probability amplitude of particle to have coordinate and momentum around $x, p$, and $\phi_p(x) = \exp(\frac{i}{\hbar}(px-Et))$, where $E = p^2/2m + U$, is the fast oscillating part. We seek an equation for the smooth envelope. We start with the Schrödinger equation in the vicinity of an external field, that is $U(x) = 0$. For time derivative of $\Psi$-function we have
\begin{equation}
i \hbar \frac{\partial}{\partial t} \Psi = i\hbar \dot A \phi_p(x) + E\Psi,
\label{time}
\end{equation}
For space derivative we have
\begin{equation}
- \frac{\hbar^2}{2m} \frac{\partial^2}{\partial x^2 }\Psi = 
\left( - \frac{\hbar^2}{2m} A''  - i\hbar \frac{p}{m} A' + \frac{p^2}{2m} A \right) \phi_p(x).
\label{space}
\end{equation}
We omitted the summation over $p$, since the equations are correct for every value of $p$ separately.

We combine equations (\ref{time}, \ref{space}). The third term in brackets in Eq. \eqref{space} cancels out with $E\Psi$. Since $A$ changes slowly with coordinate, the first term in Eq. \eqref{space} is much smaller than the second term. Taking it into account, we obtain
\begin{equation}
\dot A = - \frac{p}{m} A'
\label{preliminary}
\end{equation}

We have a similar equation for $A^*$, the complex conjugate of $A$. We multiply Eq. \eqref{preliminary} by $A^*$, and we multiply the equation for $A^*$ by $A$, and add both equations. We get
\begin{equation}
\dot \rho = - \frac{p}{m}\rho'
\label{preliminary2}
\end{equation}
where $\rho$ is the probability density of a particle to have coordinate and momentum around $x, p$, it is expressed as $\rho = A^*A$.

We now want to add the potential term. At this point, we need to introduce wave packets into the theory. We take into account that, in general, near some given point $x_0$, the $\Psi$-function may contain many components with different values of $p$, as in equation \eqref{Psi}. We substitute the $\Psi$-function in this form into the Schrödinger equation, and we do not write the kinetic-energy term, since we took care of it earlier:
\begin{equation}
i \hbar \frac{\partial}{\partial t} \sum_p A_p(x) \phi_p(x) = U(x) \sum_p A_p(x)\phi_p(x) + \dots
\label{Shred}
\end{equation}

We want to “cut out a piece’’ of the wave function in a coordinate region near $x_0$ and a momentum region near $p_0$, so we multiply the equation by $\phi_{p_0}(x)$ and integrate it over the region from $x_0$ to $x_0 + \Delta x$. For this method to be applicable, we need a scale $\Delta x$ that is large compared to the wavelengths of the fast-oscillating part of the $\Psi$-function, but small compared to the characteristic scale of variation of the amplitude.

 We notice that 
\begin{align}
\int \limits_{x_0}^{x_0 + \Delta x} \phi_{p_0}^*(x) A_p(x) \phi_p(x) \approx 
A_p(x_0) \int \limits_{x_0}^{x_0 + \Delta x} \phi_{p_0}^*(x) \phi_p(x) = 
- \frac{i\hbar}{p - p_0}\left(e^{\frac{i}{\hbar}(p - p_0) \Delta x} - 1 \right) A_p(x_0).
\label{smoothing}
\end{align}
The expression inside the brackets, as a function of $p$, is large in the vicinity of $p_0$ and small away from $p_0$. If we consider $A_p$ to be a slowly varying function of $p$, its value in the vicinity of $p_0$ can be approximated as a constant. Thus, we can put
\begin{equation}
\int_{x_0}^{x_0 + \Delta x} \phi_{p_0}^*(x) \sum_p A_p(x)\phi_p(x) \approx  \int dp \delta(p - p_0) A(x, p) = A(x_0, p_0).
\label{delta}
\end{equation}
Here we assumed that $\Delta x$ is large on the scale of the fast-varying part of the $\psi$-function $\phi_p(x)$, so the integral over the range $\Delta x$ behaves approximately as an integral over an infinite range. Here $A(x_0, p_0)$ is now a smooth function of two variables.

That is our expression for the left side of the equation. For the right side, we consider the potential $U(x)$ to be a smooth function, so we expand it into a Taylor series near $x_0$ and keep only the first term
\begin{equation}
\int_{x_0}^{x_0 + \Delta x}  \phi_{p_0}^*(x) A_p(x) U(x)\phi_p(x) = \int_{x_0}^{x_0 + \Delta x}  \phi_{p_0}^*(x) A_p(x)\left(U(x_0) + \frac{\partial U}{\partial x} x \right) \phi_p(x).
\label{Taylor}
\end{equation}
The part with $U(x_0)$ is cancelled with the energy term on the left side: $E = p^2/2 + U(x_0)$. For the term with derivative over $x$, after we have taken $\frac{\partial U}{\partial x}$ at $x_0$ outside of the integration limits, we have
\begin{align}
 \int_{x_0}^{x_0 + \Delta x}dx \phi_{p_0}^*(x) \left( \sum_p  A_p(x) x \phi_p(x) \right) 
= \int_{x_0}^{x_0 + \Delta x}dx \phi_{p_0}^*(x) \left( \sum_p A_p(x) \left[  -i \hbar \frac{\partial}{\partial p}\phi_p(x) \right] \right) = \nonumber \\
 -i \hbar \sum_p  A_p(x) \frac{\partial}{\partial p}  \left(\int_{x_0}^{x_0 + \Delta x} dx  \phi_{p_0}^*(x)   \phi_p(x) \right) =  -i \hbar \Delta x \int_p \frac{dp}{2\pi\hbar}  \frac{\partial}{\partial p} \left( \frac{2\pi\hbar}{ \Delta x}\delta(p - p_0) \right) A(x, p) = i \hbar \frac{\partial}{\partial p_0}A(x_0, p_0).
\label{derivative}
\end{align}
Here, the multiplier $\Delta x/(2\pi\hbar)$ appears because of the transition from summation to integration. This derivation relies on the interchangeability between differentiation and taking the weak limit (delta function) \cite{Zemanian2011}.

Instead of taking the derivative of a delta function, we can use a longer and less rigorous, yet more demonstrative and intuitive approach. For this, we employ a technique developed in \cite{Me24NCevo}. We divide the expression preceding $A$ on the right side of equation \eqref{smoothing} by $\Delta x$. We write
\begin{align}
- \frac{i\hbar}{\Delta x(p - p_0)}\left(e^{\frac{i}{\hbar}(p - p_0) \Delta x} - 1 \right) = \chi_{\Delta p}(p_0),
\end{align}
and observe that
\begin{align}
\Delta p \Delta x \gg \hbar =>\chi_{\Delta p}(p_0) \approx 0, \nonumber \\
\Delta p \Delta x \ll \hbar => \chi_{\Delta p}(p_0) \approx 1.
\label{cases1}
\end{align}
Now to simplify the expression, instead of $\chi_{\Delta p}(p_0)$ we introduce an approximation $\overline \chi_{\Delta p}(p_0)$ which is an indicator function of a segment of momenta in the icinity of $p_0$
\begin{align}
\Delta p \Delta x > &\pi\hbar  => \overline \chi_{\Delta p}(p_0) = 0 \nonumber \\
\Delta p \Delta x < &\pi\hbar  => \overline \chi_{\Delta p}(p_0) = 1.
\label{cases2}
\end{align}
It is a simultaneous definition of the indicator function in momentum space and of the width of the step $\Delta p$. This exchange is possible because both $\chi_{\Delta p}$ and $\overline\chi_{\Delta p}$ form a delta sequence and weakly converge to the delta function in the limit $\Delta p \to 0$.

For the first part of the derivation \eqref{delta}, this does not introduce anything new. We substitute it into the second expression from \eqref{derivative} and obtain
\begin{align}
 \int_{x_0}^{x_0 + \Delta x} \phi_{p_0}^*(x) \sum_p A_p(x) x \phi_p(x) =
 -i \hbar  \sum_p \frac{\partial}{\partial p} \left(\int_{x_0}^{x_0 + \Delta x} \phi_{p_0}^*(x) \phi_p(x) \right) A_p(x) =
 -i \hbar  \sum_p \Delta x \frac{\partial}{\partial p} \overline \chi_{\Delta p}(p_0) A_p(x) =  \nonumber \\ 
=  -i \hbar \int_p dp \frac{1}{2 \Delta p}\left( -\delta (p_0+\Delta p) + \delta (p_0 - \Delta p) \right) A(p, x_0) =  
 i \hbar \frac{A(p_0+\Delta p) - A(p_0 - \Delta p)}{2 \Delta p} =  i \hbar \frac{\partial}{\partial p_0}A(x_0, p_0),
\label{derivative2}
\end{align}
where we have used the formula for the derivative of a step function, and the last equality is the definition of the smoothed derivative. Thus, we arrive at the same conclusion in two slightly different ways.

Let us reflect on what we have achieved so far. We started with the $\psi$-function, which was a function of one variable, the coordinate, while momentum was represented as an operator. In the vicinity of external fields, momentum $p$ was an index labeling the states of the system. We partitioned the $\psi$-function into two parts -- the fast-oscillating part and a smooth envelope -- and derived an equation for the latter. It turns out that it is a function of two variables, coordinate and momentum. When we consider the equation for the smooth envelope, momentum is no longer an operator but a variable, as in classical mechanics!

We proceed to obtain $\rho$ instead of $A$ in the same manner as before, and combine it with Eq..\eqref{preliminary} to get
\begin{equation}
\dot \rho = - \frac{p}{m} \frac{\partial \rho}{\partial x} + \frac{\partial U}{\partial x} \frac{\partial \rho }{\partial p}
\label{final}
\end{equation}

If we introduce a Hamiltonian function (not operator!) $H = p^2/2m + U(x)$, we obtain
\begin{equation}
\dot \rho + \frac{\partial H}{\partial p} \frac{\partial \rho }{\partial x} - \frac{\partial H}{\partial x} \frac{\partial \rho }{\partial p} = 0,
\label{1DLiuv}
\end{equation}
the one-dimensional Liouville  equation.

We can derive Hamiltonian mechanics from this point, as a somewhat inverse Liouville's theorem. We use a method of characteristics \cite{ArnoldODE} to get a system of ordinary differential equations equivalent to this partial differential equation. We get
\begin{align}
\frac{d x}{d s} = \frac{\partial H}{\partial p} \nonumber \\
\frac{d p}{d s} = -\frac{\partial H}{\partial x}  \nonumber \\
\frac{dt}{d s} = 1  \nonumber \\
\frac{d \rho}{d s} = 0.
\label{Hamilton}
\end{align}
From the third equation we obtain $dt = ds$, so the first two equations reduce to Hamilton’s equations, while the fourth equation becomes the equation expressing conservation of probability.

While this result is trivial, especially when Liouville’s theorem is formulated in the mathematical language of Hamiltonian vector fields \cite{ArnoldMech}, we nevertheless include this derivation. It seems worthwhile, since in the physical literature authors are typically interested in deducing Liouville’s equation from Hamilton’s equations, and not the other way around.

The essence of the method of characteristics is the transition from a field description, which assigns to each point in space–time the value of the field, to a particle description, which assigns to every particle a trajectory in space–time. If we substitute the delta function in coordinate and momenta into the Liouville equation, it will continue to propagate as a delta function along the characteristic curve, and the equation of this characteristic curve is Hamilton’s equation.

In quantum mechanics, this situation corresponds to the motion of the center of coordinates and momenta of a sufficiently localized wave packet. However, Liouville’s equation admits arbitrary initial conditions, not just a delta function. In this sense, the presented theory generalizes the standard wave-packet description.

\section{Many-particle equation}

Now we want to consider the many-particle case. Let us consider the case of fermionic particles, so the fast oscillating part of the psi function can be expressed as a Slater determinant:
\begin{equation}
\psi_\mathbf{p}(\mathbf{x}) = \frac{1}{n!} \mathrm{Det} \begin{pmatrix} \phi_1(x_1) & \phi_1(x_2) & ... & \phi_1(x_n) \\ 
\phi_2(x_1) & \phi_2(x_2) & ... & \phi_2(x_n) \\
... & ... & ... & ...\\
\phi_n(x_1) & \phi_n(x_2) & ... & \phi_n(x_n) \end{pmatrix},
\label{determinant}
\end{equation} 
where  $\mathbf{x}$ stands for all the set $x_1,.. x_n$ and $\mathbf{p}$ for all the set $p_1,.. p_n$ and $\phi_k(x_l) = \exp(\frac{i}{\hbar}p_k x_l)$.

 The full wave function that contain the fast oscillating part $\psi_ \mathbf{p}(\mathbf{x})$ and the smooth envelope $A(\mathbf{x},  \mathbf{p})$ is expressed as
\begin{equation}
\Psi (\mathbf{x}) = \sum_\mathbf{p} A(\mathbf{x}, \mathbf{p}) \psi_\mathbf{p}(\mathbf{x}),
\label{PsiFull}
\end{equation}

The kinetic energy term can be expressed as 
\begin{equation}
- \frac{\hbar^2}{2m} \left( \frac{\partial^2}{\partial x_1^2 } + \frac{\partial^2}{\partial x_2^2 } + 
... + \frac{\partial^2}{\partial x_n^2} \right).
\label{KinOp}
\end{equation}

When we apply the kinetic energy operator \eqref{KinOp} to the $\psi$-function in the form (\ref{determinant}, \ref{PsiFull}), we obtain three types of summands, as in expression \eqref{space}. In the summands of the first type, the derivatives are applied twice to the slowly varying amplitude $A$, so the result is negligible. If the derivative is applied twice to the $\psi$ \eqref{determinant}, we obtain a contribution to the total kinetic energy of the system. The sum of such contributions is the total kinetic energy, and it cancels out with the time derivative of $\psi$, as in the one-particle case (\ref{time}, \ref{space}). Thus, the only summands that we should take care of are those in which one derivative is applied to $A$ and one is applied to $\psi$. The summands of this type have the form
\begin{equation}
\frac{i \hbar}{m} \frac{\partial A}{\partial x_l}
  \frac{1}{n!} \mathrm{Det} \begin{pmatrix} \phi_1(x_1)  & ... & p_1 \phi_1(x_l) & ... & \phi_1(x_n) \\ 
 \phi_2(x_1)  & ... & p_2 \phi_2(x_l) & ... & \phi_2(x_n) \\
... & ... & ... & ... & ... \\
 \phi_n(x_1)  & ... & p_n \phi_n(x_l) & ... & \phi_n(x_n) \end{pmatrix}.
\label{Summand}
\end{equation} 
Since the full fermionic psi function \eqref{PsiFull} is antisymmetric with respect to the interchange of the coordinates and the oscillating part is antisymmetric (as a property of the determinant), the amplitude should be symmetric. Hence it follows
\begin{equation}
\frac{\partial A}{\partial x_l} =  \frac{\partial A}{\partial x_k}.
\label{Symmetry}
\end{equation}
for all pairs of indeces $k, l$. 

We expand the total sum of all summands of the form \eqref{Summand}, and we expand each determinant along the column with impulses. We denote by $M_{ij}$ the determinant of the minor obtained from \eqref{determinant} after erasing the $i$-th row and the $j$-th column. We use the Laplace rule for the signs in the expansion and obtain:
\begin{align}
& \frac{\partial A}{\partial x_1} \left(p_1 \phi_1(x_1) M_{11} - p_2 \phi_2(x_1) M_{21} + ... + (-1)^{n+1} p_n \phi_n(x_1) M_{n1} \right) + \nonumber \\
& \frac{\partial A}{\partial x_2} \left(- p_1 \phi_1(x_2) M_{12} + p_2 \phi_2(x_2) M_{22} + ... + (-1)^{n+2} p_n \phi_n(x_2) M_{n2} \right) + ...
\end{align} 
We expand the brackets and rearrange the partial derivatives with the help of expression \eqref{Symmetry} in such a way that the index of the derivative coincides with the index of the momentum:
\begin{align}
& \frac{\partial A}{\partial x_1} p_1 \phi_1(x_1) M_{11} - \frac{\partial A}{\partial x_2} p_2 \phi_2(x_1) M_{21} + ...  \nonumber \\
& - \frac{\partial A}{\partial x_1} p_1 \phi_1(x_2) M_{12} + \frac{\partial A}{\partial x_2} p_2 \phi_2(x_2) M_{22} + ... 
\end{align} 
We finally group members with the same index of momentum and derivative
\begin{align}
& \frac{\partial A}{\partial x_1}p_1 \left( \phi_1(x_1) M_{11} - \phi_1(x_2) M_{12} + ... + (-1)^{n+1}\phi_1(x_n) M_{1n} \right) + \nonumber \\
& \frac{\partial A}{\partial x_2}p_2 \left( - \phi_2(x_1) M_{21} + \phi_2(x_2) M_{22} + ... + (-1)^{n+2}\phi_2(x_n) M_{2n} \right) + ...
\end{align} 
The expressions in the brackets are nothing but the expansions of the determinant over the rows. The expression in the first line of the equation is the expansion over the first row, the expansion in the second line is the expansion over the second row, and so on. However, every such expansion is just the same determinant $\psi_ \mathbf{p}(\mathbf{x})$ \eqref{determinant}. Thus, we have
\begin{align}
 \left(\frac{\partial A}{\partial x_1}p_1 + \frac{\partial A}{\partial x_2}p_2 + ... \right) \psi_ \mathbf{p}(\mathbf{x}).
\end{align} 
We note that the same expression would arise for bosons, except that in this case the determinant would be replaced by the permanent in the expression for $\psi_\mathbf{p}(\mathbf{x})$ \eqref{determinant}. The permanent is analogous to the determinant, but without the alternating signs between summands. All that we need to make an analogous derivation is the possibility of expanding the permanent by minors along a row or a column, a property that it shares with the determinant.
 
Since the total $\psi$-function should be symmetric and the permanent is also symmetric, the amplitude should be symmetric as well, so that expression \eqref{Symmetry} holds. Otherwise, the derivation is the same. We note that in books on physical kinetics \cite{Harris2004}, the distribution function in Liouville's equation is assumed to be symmetric. Here, we have justified this assumption, both for fermions and bosons.

For the momentum part of our derivation, we introduce a region of space bounded between $x_{01}$ and $x_{01} + \Delta x_{01}$, $x_{02}$ and $x_{02} + \Delta x_{02}$, and so forth. Furthermore, when we use integration over coordinates, we mean integration over this region. We expand the potential function inside this region as
\begin{equation}
U(\mathbf{x}) = U(\mathbf{x}_0) + \frac{\partial U}{\partial x_1} x_1 + ... + \frac{\partial U}{\partial x_n} x_n.
\label{TaylorND}
\end{equation}
We notice, that 
\begin{equation}
x_l \psi_\mathbf{p}(\mathbf{x}) =
 \mathrm{Det} \begin{pmatrix}
 \phi_1(x_1)  & ... & \frac{\partial}{\partial p_1} \phi_1(x_l) & ... & \phi_1(x_n) \\ 
 \phi_2(x_1)  & ... & \frac{\partial}{\partial p_2} \phi_2(x_l) & ... & \phi_2(x_n) \\
... & ... & ... & ... & ... \\
 \phi_n(x_1)  & ... & \frac{\partial}{\partial p_n} \phi_n(x_l) & ... & \phi_n(x_n) \end{pmatrix} =
\sum_k \frac{\partial}{\partial p_k} \phi_k (x_l) M_{kl}.
\label{MinorsX}
\end{equation} 
Also we notice that
\begin{equation}
 \int d x_l \phi_m^* (x_l) \frac{\partial}{\partial p_k}  \phi_k (x_l) = \delta_{mk}  \frac{\partial}{\partial p_k} \delta(p_k - p_{0k}).
\label{IntDeltaN}
\end{equation}

We multiply the Schrödinger equation over $\psi_\mathbf{p_0}(\mathbf{x})^\dag$, integrate over a region of space, substitute the multidimensional Taylor expansion for potential, and get, combining everything above:
\begin{align}
 \int d\mathbf{x} \psi_\mathbf{p_0}^\dag(\mathbf{x}) \left ( \sum_\mathbf{p} A_\mathbf{p}(\mathbf{x}) \left[ \sum_i \frac{\partial U}{\partial x_i} x_i \right] \psi_\mathbf{p}(\mathbf{x})  \right) =
 \nonumber \\ 
=  \int d\mathbf{x} \psi_\mathbf{p_0}^\dag(\mathbf{x}) \left ( \sum_\mathbf{p} A_\mathbf{p}(\mathbf{x}) \left[ \sum_i \frac{\partial U}{\partial x_i} \sum_j \frac{\partial }{\partial p_j}  \phi_j  M_{ij} (\mathbf{x}) \right] \right)
=
 \nonumber \\ 
=    \sum_\mathbf{p} A_\mathbf{p}(\mathbf{x})  \sum_i \frac{\partial U}{\partial x_i} \sum_j  \left[ \int d\mathbf{x} \psi_\mathbf{p_0}^\dag(\mathbf{x}) \frac{\partial }{\partial p_j}  \phi_j  M_{ij} (\mathbf{x}) \right] 
=
 \nonumber \\ 
=    \sum_\mathbf{p} A_\mathbf{p}(\mathbf{x})  \sum_i \frac{\partial U}{\partial x_i} \prod_{j\neq k} \delta(p_j - p_{0j})  \delta_{ik}  \frac{\partial}{\partial p_k} \delta(p_k - p_{0k})  =
 \nonumber \\ 
= \sum_i \frac{\partial U}{\partial x_i} \frac{\partial}{\partial p_i} A_{\mathbf{p}_0}(\mathbf{x}_0)  
\end{align}

Overall, we get the equation of evolution of the amplitude for arbitrary many degrees of freedom:
\begin{equation}
\dot A + \sum_{i=1}^n \left( \frac{\partial H}{\partial p_i} \frac{\partial A }{\partial x_i} - \frac{\partial H}{\partial x_i} \frac{\partial A }{\partial p_i} \right) = 0.
\label{PreLiouville}
\end{equation}
We see that this equation can be turned into the Liouville equation for the probability density, with the same manipulation we used before (\ref{preliminary}, \ref{preliminary2}).

\section{Applications to physical kinetics}
In this section, we combine the results obtained in Sections II and III with the results of our previous papers to derive the main equations of physical kinetics.

If we exclude interparticle interactions and allow the potential term to describe some smooth external potential, then averaging Liouville’s equation indeed yields the non-collision part of the Boltzmann equation:
\begin{equation}
\dot f(x, p) + \left( \frac{\partial H}{\partial p} \frac{\partial f }{\partial x} - \frac{\partial H}{\partial x} \frac{\partial f }{\partial p} \right) = 0
\label{Boltzmann}
\end{equation}

However, we could have arrived at this result much more simply, without all the complicated manipulations with the many-particle $\psi$-function of Section III. In the spirit of Ref. \cite{Me24transmission}, we can express the $\psi$-function as
\begin{equation}
\hat \Psi (x) = \sum_p \hat a_p(x) \phi_p(x).
\label{PsiHat}
\end{equation}
where $\hat a_p(x)$ represents the destruction operator of a particle, no matter bosonic or fermionic. If we perform all the manipulations of Section II with expression \eqref{PsiHat} we arrive at 
\begin{equation}
\dot{\hat a}_p + \frac{\partial H}{\partial p} \frac{\partial \hat a_p }{\partial x} - \frac{\partial H}{\partial x} \frac{\partial \hat a_p }{\partial p} = 0.
\label{BoltzmannHat}
\end{equation}
 
Further reasoning is similar to what we did for the transition between Eqs. \eqref{preliminary} and \eqref{preliminary2}. We multiply equation \eqref{BoltzmannHat} by $\hat a_p^\dag$. We also have an equation analogous to \eqref{BoltzmannHat}, but for $\hat a_p^\dag$, and we multiply this equation by $\hat a_p$. We sum these two equations and obtain an equation that is very similar to the Boltzmann equation, but for the number-of-particles operator $\hat n_p = \hat a_p^\dag \hat a_p$. We put bra and ket vectors on both sides, and since $\langle S|\hat n_p | S \rangle = n_p$, the mean number of particles with momenta $p$ in the state $| S \rangle$, we arrive at the Boltzmann equation.

 In Ref. \cite{Me24transmission} we used more general formula for the mean number of particles
\begin{align}
\langle S| \hat a_p^\dag \hat a_{p'} | S \rangle = n_p \delta_{pp'},
\label{2Kronecker}
\end{align} 
but we were not in a position to derive it accurately. Let us do it here, for further use.

Formula \eqref{2Kronecker} is conditional. It is valid only when the system is in a non-coherent state. This means that, at a given point, the probability amplitudes of different microstates are not phase-correlated with each other. The exact physical conditions that allow us to use this physical assumption, as well as its consequences, are described in \cite{Me24NCevo} in great detail.

To explain this briefly, we recall that we are studying an extension of the wave-packet theory. Therefore, the state of the system consists of a sheaf of frequencies. When two states interact with each other, the interaction of each pair of single-frequency components is described, taking into account the phase difference between such components. However, the phase difference is a function of frequency. And for different frequencies, the phase difference is different. When calculating the mean value over all frequencies, if no phase correlation is present, we obtain values as if the states did not have amplitudes with phases, but only probabilities.

The standard expression for the mean value of some physical quantity in quantum mechanics is:
\begin{equation}
\langle O \rangle = \sum_{k,l} A_k^* O_{k l} A_l.
\label{MeanQ}
\end{equation}
As we have shown in \cite{Me24NCevo}, if states $k, l$ are not phase-correlated, it is equivalent to introducing a delta-symbol into the summation
\begin{equation}
\langle O \rangle = \sum_{k,l} A_k^* O_{kl} A_\beta \delta_{kl} = \sum_{k} |A_k|^2 O_{kk} =  \sum_{k} P_k O_k,
\label{MeanS}
\end{equation}
where $P_k$ is the probability for the system to be in state $k$ and $O_k$ is the value of the physical quantity in this state (short for $O_{kk}$). Hence, for non-coherent systems, the expression for the mean value in quantum mechanics reduces to the standard expression for the mean value in statistical physics.

We can use this to derive formula \eqref{2Kronecker}. The case when $p = p'$ is obvious, and we have already used it earlier. Let us consider the case $p \neq p'$. We substitute $k = ij$, $k' = i'j'$ into formula \eqref{MeanS}, where $i, i'$ are the numbers of particles in the microstate with momentum $p$, and $j, j'$ in $p'$, and obtain
\begin{align}
\langle S| \hat a_p^\dag \hat a_{p'} | S \rangle = \sum_{ij,i'j'} A_{ij}^* \langle i, j| \hat a_p^\dag \hat a_{p'} | i', j' \rangle A_{i'j'}^* \delta_{ii'} \delta_{jj'}.
\end{align} 
We see that delta symbols are non-zero only in case $i=i', j=j'$, while the expression with operators equals zero in this case. So the formula \eqref{2Kronecker} is proven.

Let us use this formula to find an expression for the current density. We write the current operator standardly
\begin{equation}
\hat{\vec J} = \frac{\hbar e}{2mi}\left( \hat \Psi^\dag \vec \nabla \hat \Psi - \hat \Psi \vec \nabla \hat \Psi^\dag \right),
\label{Current}
\end{equation}
and substitute $\hat \Psi$ in the form \eqref{PsiHat} into it. Neglecting the small quantity $ \vec\nabla \hat a_p$, we get 
\begin{equation}
\hat{ \vec J} = \frac{\hbar e}{2mi} \frac{i}{\hbar} \sum_{\vec p,\vec p'} \left(\vec p' a^\dag_{\vec p} a_{\vec p'} - \vec p a_{\vec p} a^\dag_{\vec p'}\right).
\label{Current2}
\end{equation}

As we have found \eqref{2Kronecker}, if $\vec p \neq \vec p'$ the contribution goes to zero after averaging over states, so we only keep summands with $\vec p = \vec p'$
\begin{equation}
\hat{ \vec J} = e \sum_{\vec p} \frac{\vec p}{m}\left( a^\dag_{ \vec p} a_{\vec p} + 1/2 \right).
\label{Current3}
\end{equation}
Because of symmetry, the summation of constant over $\vec p$ equals zero. After averaging over states and going from summation to integration, we get 
\begin{equation}
\vec j (x) = e \int \frac{d^3 \vec p}{(2\pi \hbar)^3} \frac{\vec p}{m} f(x),
\label{Current4}
\end{equation}
which is a standard expression for the current density in physical kinetics. We see that, first, the use of the techniques that we have introduced allows us to write the expression for the current locally, at a single point, instead of writing averages over the whole space. Second, the current is expressed as a sum of the currents produced by each particle, that is, the product of the charge and the velocity $e p/m$. This means that, under the given assumptions, the particles of the system under consideration behave as if they were classical point particles.

We have obtained the formula for the charge current density, but the derivation for any other current density is analogous. The derivation for the densities of charge, energy, and similar quantities is also analogous and even simpler, since we can start from $( \hat \Psi^\dag \hat \Psi$, which contains no derivatives.

Finally, we want to derive the full Boltzmann equation from quantum mechanics. We assume that the only smooth potential in the system is some externally applied potential $\sum U(x_i)$, and that the interparticle interactions are “sharp”, or, in other words, short-ranged.
\begin{equation}
\Psi (\mathbf{x}) = \sum_\mathbf{p} A(\mathbf{x}, \mathbf{p}, t) \psi_\mathbf{p}(\mathbf{x}, t),
\label{PsiFinall}
\end{equation}
and the dependency of $\psi$ on $t$ is given by exponent
\begin{equation}
\psi (t) = \exp \left(\frac{iHt}{\hbar} \right); H = \sum_\mathbf{p} \frac{p_i^2}{2m} + U(\mathbf{x}).
\label{Hamilton}
\end{equation}

We perform the same operation of multyplying the Schrödinger equation over $\psi_{\mathbf{p}, H}(\mathbf{x}_0, t_0)$ and integrating over space intervals $x_{01}$ and $x_{01} + \Delta x_{01}$, $x_{02}$ and $x_{02} + \Delta x_{02}$ and so forth, and a time interval $t_{0} + \Delta t$. We use the same reasoning as in Section III to treat the external potential and space dependence of amplitudes. And the same reasoning as in Ref. \cite{Me24NCevo} to treat interparticle interaction. We arrive at
\begin{equation}
\dot A_k + \sum_{i=1}^n \left( \frac{\partial H}{\partial p_i} \frac{\partial A_k }{\partial x_i} - \frac{\partial H}{\partial x_i} \frac{\partial A_k }{\partial p_i} \right) = \sum_{l} V_{kl} A_k,
\label{FinalAmps}
\end{equation}
where $V_{kl}$ are matrix elements of interparticle interactions and $V_{kk} = 0$.

We go from amplitudes to probability densities and invoke the non-coherence condition for the right side:
\begin{equation}
\dot \rho_k + \sum_{i=1}^n \left( \frac{\partial H}{\partial p_i} \frac{\partial \rho_k }{\partial x_i} - \frac{\partial H}{\partial x_i} \frac{\partial \rho_k }{\partial p_i} \right) = \sum_{l} Q_{kl} \rho_k.
\label{FinalRhos}
\end{equation}
Here $Q_{kl}$ are elements of the transition rate matrix. Transitions are between states of the system at given point and its elements are 
\begin{align}
Q_{kl} = \frac{2 \pi}{\hbar} |V_{kl}|^2 \delta(E_k - E_l), \nonumber \\
Q_{kk} = -  \frac{2 \pi}{\hbar} \sum_l |V_{kl}|^2 \delta(E_k - E_l).
\label{Fermi}
\end{align}
For non-diagonal elements, these transition rates coincide with the Fermi rule, and the diagonal values have such a value that the sum of all elements in each row equals zero.

Finally, we calculate the evolution of a mean value of the number of particles at a given point and get
\begin{equation}
\dot f(x, p) + \left( \frac{\partial H}{\partial p} \frac{\partial f }{\partial x} - \frac{\partial H}{\partial x} \frac{\partial f }{\partial p} \right) = \text{St} (f).
\label{FullBoltzmann}
\end{equation}
The right side is a collision integral. We have shown this in Ref. \cite{Me24NCevo} for a few frequently used examples. Equation \eqref{FullBoltzmann} is the well-known Boltzmann equation, which includes both the collisional and the non-collisional terms, and which we have derived from the Schrödinger equation.

\section{Summary and discussion}

It is typically thought that the Boltzmann equation is derived from classical mechanics, and it is assumed that classical particles are classical because they are considered “heavy”. Meanwhile, the Boltzmann equation is successfully used in the description of kinetic phenomena in solid states \cite{Ziman1960, LandauV10}, where the particles are quasiparticles, are described by quantum distribution functions (Bose–Einstein and Fermi–Dirac), and do not possess momentum, but rather quasimomentum. Concerning masses, electrons and holes typically have effective masses that are a fraction of the electron mass \cite{Anselm1981}, which is very “light”, and phonons are massless quasiparticles.

Yet thinking about them as point particles that are, at least partly, described by classical mechanics provides fruitful physical insights into transport phenomena in crystals. For example, the theoretical prediction of the properties of a field-effect transistor is based on the diffusion equation \cite{Anselm1981}, which more or less explicitly assumes the object of its description to behave like a classical gas. Such an application should have a proper justification.

The applicability of the mental model of a classical gas follows from the applicability of the Boltzmann equation and of classical expressions for densities and current densities. Their systematic derivation from quantum mechanics justifies the use of classical concepts for quasiparticles. A serious obstruction to this justification is the completely different mathematical methods that are used to derive the different parts of the Boltzmann equation.

The collision-integral part consists of instantaneous transitions of the system from one state to another, and the rate of such transitions, in the simplest case, is computed using the Fermi rule. The method employed by Dirac to derive the Fermi rule, which was used and extended in our previous work \cite{Me24NCevo}, was based on the smooth envelope concept. We seek the wave function in the form $\Psi = \sum A \psi$, where $A$ is an amplitude slowly varying in time and $\psi$ is the wave function of an interactionless system.

The non-collision part of the Boltzmann equation is a consequence of the Liouville equation, which can be derived from classical mechanics. And we already know that classical mechanics can be derived from quantum mechanics. However, we see that the standard methods of derivation of classical mechanics (and, consequently, of the non-collisional part of the Boltzmann equation) are completely different from the method of derivation of the collision-integral part. Our goal was to derive the Boltzmann equation as a single piece from the Schrödinger equation, and to see where the line lies that separates the first part from the second part.

To accomplish this, we derive the Liouville equation here directly from Schrödinger’s equation, using the smooth-envelope technique—the same one we previously used to describe quantum transitions (and, consequently, the collision integral). We see that fields that vary slowly, that is, with approximately the same rate as the amplitudes of the states, act on the system in such a way that it exhibits classical-like behavior, which is reversible and deterministic. By contrast, the interaction of the system with fast-oscillating fields results in behavior characterized by quantum transitions, which is probabilistic in nature and irreversible.

The belief that the Boltzmann equation is derived through classical mechanics is known to lead to a very serious contradiction: classical mechanics is time-reversible, while the Boltzmann equation is irreversible. We carefully address the emergence of irreversibility in our theory and come to the conclusion that the source of irreversibility is non-coherence. When two waves overlap, the non-coherence devours interference and produces irreversibility.

In the derivation presented in this manuscript, we have seen that the amplitudes satisfy the same equation \eqref{PreLiouville} as the probability densities, namely, the Liouville equation. This equation, by the famous Liouville theorem \cite{ArnoldMech}, conserves measure, which, when translated into simple language, means that different small parts of the envelope function, following their prescribed trajectories \eqref{Hamilton}, do not overlap. This guarantees that no interference arises, and that is why no irreversibility is present in the non-collisional part of the Boltzmann equation. This is in contrast with the collisional part, where the system can transit to one particular state from a handful of other states, and the amplitudes of these initial states interfere with each other.

This resolves the paradox of how the irreversible Boltzmann equation emerges from reversible classical mechanics. Our reasoning suggests that it does not. The non-collision part and the collision part are two different limits of quantum mechanics, which are the long-wavelength and short-wavelength limits, respectively.

Moreover, as we have noted in the introduction, the theory presented in this manuscript is an extension of the wave-packet theory. The wave packets, which represent the motion of a particle along a trajectory, are assumed to have a finite spectral width. In our theory of non-coherent evolution, which was presented earlier, we also assumed a finite spectral width of the states in order to explain why the reversible unitary evolution is replaced by irreversible stochastic evolution. We see that the two complementary parts of our theory rely on the same physical assumptions: the representability of the wave function as a composition of two parts (slowly oscillating and fast oscillating), and the finite spectral width of the states.

The conclusions of these two articles are a result of these assumptions and sufficiently rigorous mathematical manipulations. We consider the behaviour of the wave function that fulfills the assumptions, in interaction with fields in two limiting cases (low-frequency and high-frequency fields), and obtain the two parts of the Boltzmann equation.

There is an alternative theory that explains the right part of equation \eqref{FinalRhos}, which leads to the collision-integral part of the Boltzmann equation, known as quantum thermodynamics \cite{breuer2002openquant}. In this context, the right part of equation \eqref{FinalRhos} is known as the Born–Markov approximation of the Lindblad master equation \cite{lindblad1976generators}. Quantum thermodynamics has successfully shown that a system described by the master equation tends to thermal equilibrium \cite{rigol2008thermalization, deutsch2010thermodynamic, dymarsky2018subsystem}. However, applying it in the context of physical kinetics seems problematic.

Quantum thermodynamics attributes the loss of coherence in a quantum system to its interaction with the environment. To obtain the correct form of the Lindblad equation, one must assume that the heat-bath relaxation time is small compared to the relaxation time of the system \cite{breuer2002openquant}. This essentially means that there is an additional interaction that is stronger than the interaction of interest. If such an interaction exists, it should result in a reduction of the particles' free path and thus contribute to the kinetic coefficients. However, in quantum thermodynamics this does not happen: the external interaction only affects the form of the evolution equation, making it irreversible, but does not affect its parameters. Our theory is free of such a contradiction.

In \cite{Me24NCevo} we noted that our methods take into account only contributions from waves with wavelengths much smaller than $\Delta x$, the technical scale of integration (see Eq.~\eqref{smoothing} and the discussion there). Information about low-frequency components was thus lost, and we suggested that such low-frequency modes should be treated as the force term in the Boltzmann equation. Here, we have successfully provided such a description.

What is lacking in our derivation is a formal mathematical procedure for separating the potential into two parts, one of which would be “sharp” and contribute to the collision integral, while the other would be “smooth” and contribute to the force term. In Section~IV of the manuscript, we specifically required the interparticle interaction to be short-ranged, since our methods do not allow for a systematic treatment of long-range potentials. Resolution of this problem would not only extend the results of this paper to a broader class of systems, including those with long-range interactions. A systematic exclusion of small-frequency contributions from integrals arising in perturbation theory is very important in cases where such integrals are divergent at zero frequency. This type of divergence is known as an infrared divergence. A formal separation of the potential into sharp and smooth parts, of which only the sharp part contributes to perturbation-theory integrals, would provide a physically clear and mathematically rigorous procedure for eliminating infrared divergences.

\bibliography{timerev, core}
\bibliographystyle{ieeetr}

\end{document}